OpenStudy (anonymous):

Please Help! I got cot2a = -7/24. How do you make it into cos2a=-7/25!? Im so lost!

OpenStudy (anonymous):

i really yneed help!!

OpenStudy (anonymous):

ok we can do this although it is not necessarily true, it could be the case that \(\cos(2a)=\frac{7}{25}\)

OpenStudy (unklerhaukus):

 \[\cos\theta=\frac{\text {adj}}{\text {hypo}}\] \[\cot\theta=\frac{\text {adj}}{\text {opp}}\]